# Orientable and negative orientable sequences

### v1

@@ 4.1 Preliminaries
Suppose that $q>2$, $S$ is a $\mathcal{NOS}_{q}(n)$ of period $m$ containing all the $i$-uniform $n$-tuples ($i,i,i\ldots,i$) for every $i$ satisfying $21\leq i<q/$ 2, and $S^{\prime}$ is derived from $S$ using Construction 4.3. Then $S^{\prime}$ is a $\mathcal{NOS}_{q}(n)$ with $w_{q}(S^{\prime})$ a unit in $\mathbb{Z}_{q}$, and period at least $m-1$ ($q\neq 6$) or at least $m-2$ ($q=6$).
@@ 4.2 Orientable sequences for $n=3$
We can how give a method for constructing near-optimal orientable sequences for $n=3$ and any $q>2$.
Construction 4.5
Suppose $q>2$ and let $S$ be a maximal $\mathcal{NOS}_{q}(2)$ constructed using the method described in Section 3.2. Let $S^{\prime}$ be derived from $S$ using Construction 4.3 (observing that $S$ satisfies the conditions of the construction by Note 3.12). Then let $T$ be derived from $S^{\prime}$ using Theorem 4.1.
@@
Suppose $q>2$ and suppose that $T$ is constructed from $S$ using Construction 4.5. Then $T$ is an $\mathcal{OS}_{q}(3)$ of period at least:
$q(\frac{q(q-1)}{2}-1)~{}~{}$ if $q$ is odd;
$q(\frac{q(q-1)}{2}-2)~{}~{}$ if $q$ is even ($q\not=6$);
$q(\frac{q(q-1)}{2}-3)~{}~{}$ if $q=6$.
@@ 4.3 Orientable sequences for general $n$
Construction 4.7
Suppose $q>2$ and $n>2$, and let $S$ be a $\mathcal{NOS}_{q}(n-1)$ constructed using the method described in Section 3.3. Let $S^{\prime}$ be derived from $S$ using Construction 4.3 (observing that $S$ satisfies the conditions of the construction as noted above). Then let $T$ be derived from $S^{\prime}$ using Result 4.1.
@@
Suppose $n>2$, $q>2$, and $T$ is derived using Construction 4.7. Then $T$ is an $\mathcal{OS}_{q}(n)$ of period at least:
@@
    By Lemma 3.15 there exists a $\mathcal{NOS}_{q}(n-1)$ $S$ of period $(q^{n-1}-r_{n,q-1})/2$. By Lemma 4.4, constructing $S^{\prime}$ from $S$ using Construction 4.3 will yield a $\mathcal{NOS}_{q}(n-1)$ of period:
@@
    The result follows from Result 4.1.
By Note 3.18, these sequences have period close to the optimal.
@@ 4.4 Examples
As noted above, the $\mathcal{NOS}_{3}(2)$ $S_{3,2}=[011]$ has $w_{3}(S_{3,2})=2$, a unit in $\mathbb{Z}_{3}$ and hence can be used directly in Result 4.1. Applying $D^{-1}$ to $S_{3,2}$ results in the $\mathcal{OS}_{3}(3)$ with ring sequence $[001220112]$, of period 9. This is optimal (see [2, Table 1]).
The $\mathcal{NOS}_{3}(3)$ $S_{3,3}=[0100112111]$ has $w_{3}(S_{3,3})=2$, which is again a unit in $\mathbb{Z}_{3}$ and hence it too can be used directly in Result 4.1. Applying $D^{-1}$ to $S_{3,3}$ results in the $\mathcal{OS}_{3}(4)$ with ring sequence
@@
of period 27. This is six less then the bound of [2, Theorem 4.11].
Finally, the $\mathcal{NOS}_{4}(3)$ $S_{4,3}$ has $w_{4}(S_{4,3})=0$, which is not a unit in $\mathbb{Z}_{4}$. To use it in Result 4.1 we need to remove a single ‘1’ from the 1-uniform 3-tuple (111), resulting in the $\mathcal{NOS}_{4}(3)$
@@
5 Further large-period orientable sequences
In this final main section we describe how to construct orientable sequences of large period using a second approach described by Alhakim et al. [2].
@@ 5.2 Recursively constructing orientable sequences
[…]
$(q-1)(q-2)/2-2~{}~{}~{}\mbox{if $q=6$},$
@@
[…]
$((q-1)^{3}-(3q^{2}-6q+4)/4)/2-1$ if q is even ($~{}~{}~{}\mbox{if $q$ is even ($q\neq 6$)},$
$((q-1)^{3}-(3q^{2}-6q+4)/4)/2-2$ if $~{}~{}~{}\mbox{if $q=6$}.$
@@
    First observe that if $S$ is a negative orientable sequence obtained using the method underlying Lemma 3.21 then it is good by definition (since it contains no zeros). It also clearly contains all the $i$-uniform $n$-tuples $(i,i,\ldots,i)$ for every $i$ satisfying $21\leq i<q/$ 2. If it has period $m$ then, using Construction 4.3, we can construct a good $\mathcal{NOS}_{q}(n)$ $S^{\prime}$ where $w(S^{\prime})$ is a unit in $\mathbb{Z}_{q}$, with period at least $m-1$ (or $m-2$ if $q=6$).
@@
This has asymptotically optimal period as $q\rightarrow\infty$.
@@ 6 Concluding remarks
In this paper we used two of the approaches proposed in [2] to generate orientable sequences with large period. Whilst the second approach yields sequences with somewhat shorter periods, in practical applications it may be that its recursive (and explicit) approach is more convenient.
It remains to consider how to devise potential input sequences for the other methods of construction proposed there.

### v5

@@ 4.2 Construction methods
Suppose that $q>2$, $S$ is a $\mathcal{NOS}_{q}(n)$ of period $m$ containing all the $i$-uniform $n$-tuples ($i,i,i\ldots,i$) for every $i$ satisfying $21\leq i<q/$ 2, and $S^{\prime}$ is derived from $S$ using Construction 4.4. Then $S^{\prime}$ is a $\mathcal{NOS}_{q}(n)$ with $w_{q}(S^{\prime})$ a unit in $\mathbb{Z}_{q}$, and period at least $m-1$ ($q\neq 6$) or at least $m-2$ ($q=6$).
@@ 4.3 Orientable sequences for $n=3$
We can now give a method for constructing near-optimal orientable sequences for $n=3$ and any $q>2$. These have period asymptotical of the same order as the maximum possible as $q\ \rightarrow\infty$.
Construction 4.6
Suppose $q>2$ and let $S$ be a maximal $\mathcal{NOS}_{q}(2)$ constructed using the method described in Section 3.2. Let $S^{\prime}$ be derived from $S$ using Construction 4.4 (observing that $S$ satisfies the conditions of the construction by Note 3.12). Then let $T$ be derived from $S^{\prime}$ using Theorem 4.2.
@@
Suppose $q>2$ and suppose that $T$ is constructed from $S$ using Construction 4.6. Then $T$ is an $\mathcal{OS}_{q}(3)$ of period at least:
$q\left(\frac{q(q-1)}{2}-1\right)~{}~{}$ if $q$ is odd;
$q\left(\frac{q(q-1)}{2}-2\right)~{}~{}$ if $q$ is even ($q\not=6$);
$q\left(\frac{q(q-1)}{2}-3\right)~{}~{}$ if $q=6$.
@@
Periods of the orientable sequences obtained using Theorem 4.7 are tabulated for small $q$ in Table 13. The value of the upper bound given in [2] is given in brackets for comparison purposes.
Table 13: $\mathcal{OS}_{q}(3)$ periods (and bounds)
$q=3$ $q=4$ $q=5$ $q=6$ $q=7$ $q=8$
6 (9) 16 (22) 45 (50) 72 (87) 140 (147) 208 (220)
@@ 4.4 Orientable sequences for general $n$
Construction 4.8
Suppose $q>2$ and $n>2$, and let $S$ be a $\mathcal{NOS}_{q}(n-1)$ constructed using the method described in Section 3.3. Let $S^{\prime}$ be derived from $S$ using Construction 4.4 (observing that $S$ satisfies the conditions of the construction as noted above). Then let $T$ be derived from $S^{\prime}$ using Result 4.2.
@@
Suppose $n>2$, $q>2$, and $T$ is derived using Construction 4.8. Then $T$ is an $\mathcal{OS}_{q}(n)$ of period at least:
@@
    By Lemma 3.15 there exists a $\mathcal{NOS}_{q}(n-1)$ $S$ of period $(q^{n-1}-r_{n,q-1})/2$. By Lemma 4.5, constructing $S^{\prime}$ from $S$ using Construction 4.4 will yield a $\mathcal{NOS}_{q}(n-1)$ of period:
@@
    The result follows from Result 4.2.
By Note 3.18, these sequences have period of the same order of magnitude as the upper bound. Periods of the orientable sequences obtained using Theorem 4.9 are tabulated for small $q$ in Table 14. As previously, the value of the upper bound given in [2] is given in brackets for comparison purposes.
Table 14: $\mathcal{OS}_{q}(n)$ periods (and bounds)
$n$ $q=3$ $q=4$ $q=5$ $q=6$ $q=7$ $q=8$
3 6 16 45 72 140 208
(9) (22) (50) (87) (147) (220)
4 27 84 275 522 1127 1792
(33) (118) (290) (627) (1155) (2012)
5 90 368 1385 3288 7756 14672
(105) (478) (1490) (3777) (8211) (16124)
6 285 1540 7155 20160 56049 118856
(336) (2014) (7680) (23217) (58464) (130812)
7 879 6340 35805 122634 388619 959152
(1032) (8062) (38640) (139317) (410256) (1046524)
8 2688 25900 181095 743358 2757930 7724544
(3189) (32638) (194630) (839157) (2879835) (8386556)
@@ 4.5 Examples
As noted above, the $\mathcal{NOS}_{3}(2)$ $S_{3,2}=[011]$ has $w_{3}(S_{3,2})=2$, a unit in $\mathbb{Z}_{3}$ and hence can be used directly in Result 4.2. Applying $D^{-1}$ to $S_{3,2}$ results in the $\mathcal{OS}_{3}(3)$ with ring sequence $[001220112]$, of period 9. This is optimal (see [2, Table 1]).
The $\mathcal{NOS}_{3}(3)$ $S_{3,3}=[0100112111]$ has $w_{3}(S_{3,3})=2$, which is again a unit in $\mathbb{Z}_{3}$ and hence it too can be used directly in Result 4.2. Applying $D^{-1}$ to $S_{3,3}$ results in the $\mathcal{OS}_{3}(4)$ with ring sequence
@@
of period 27. This is six less than the bound of [2, Theorem 4.11].
Finally, the $\mathcal{NOS}_{4}(3)$ $S_{4,3}$ has $w_{4}(S_{4,3})=0$, which is not a unit in $\mathbb{Z}_{4}$. To use it in Result 4.2 we need to remove a single ‘1’ from the 1-uniform 3-tuple (111), resulting in the $\mathcal{NOS}_{4}(3)$
@@
5 More orientable sequences
In this final main section we describe how to construct orientable sequences using a second approach described by Alhakim et al. [2]. These have period of the same order of magnitude as the upper bound.
@@ 5.2 Recursively constructing orientable sequences
$m_{2}=\begin{cases}(q-1)(q-2)/2-1,&\text{if~{}}q\neq 6\\ (q-1)(q-2)/2-2,&\text{if~{}}q=6\end{cases}$
@@
$m_{3}=\begin{cases}(q-1)^{3}/2-1,&\text{if $q$ is odd}\\ ((q-1)^{3}-(3q^{2}-6q+4)/4)/2-1,&\text{if $q$ is even ($q\neq 6$)}\\ ((q-1)^{3}-(3q^{2}-6q+4)/4)/2-2,&\text{if $q=6$}.\end{cases}$
@@
    First observe that if $S$ is a negative orientable sequence obtained using the method underlying Lemma 3.21 then it is good by definition (since it contains no zeros). It also clearly contains all the $i$-uniform $n$-tuples $(i,i,\ldots,i)$ for every $i$ satisfying $21\leq i<q/$ 2. If it has period $m$ then, using Construction 4.4, we can construct a good $\mathcal{NOS}_{q}(n)$ $S^{\prime}$ where $w(S^{\prime})$ is a unit in $\mathbb{Z}_{q}$, with period at least $m-1$ (or $m-2$ if $q=6$).
@@
For each $s$ these have period of the same order of magnitude as the upper bound.
@@ 6 Concluding remarks
In this paper we used two of the approaches proposed in [2] to generate orientable sequences with period of the same order of magnitude as the upper bound. Whilst the second approach yields sequences with somewhat shorter periods, in practical applications it may be that its recursive (and explicit) approach is more convenient.
There remain a variety of directions for future research. The constructions in this paper require an exponential amount of space with respect to $n$. They either require storage of a graph of exponential size to use an Eulerian circuit algorithm, or storage of the sequence of a smaller order to use Lempel’s homomorphism. It would be desirable to be able to generate negative orientable sequences more space-efficiently. Another potential future research direction would be to generate longer negative orientable sequences by incorporating some words with pseudoweight equal to $nq/2$. Finally, it remains to consider how to devise potential input sequences for the other methods of construction proposed in [2].